OpenStudy (anonymous):

In a direct variation, y is negative two when x is 4. What is y when x equals negative fourteen?

OpenStudy (johnweldon1993):

Direct variation means in the form \[\large y = kx\] So when x = 4...y = -2 , pug those in... \[\large -2 = 4k\] Divide both sides by 4 to find out what k is... \[\large k = -\frac{1}{2}\] Now that we know k = -1/2 we can solve the second part of the question... when x = -14 y = ? \[\large y = kx\] \[\large y = -\frac{1}{2}x\] When x = -14... \[\large y = -\frac{1}{2}\times -14\] \[\large y = \frac{14}{2}\] \[\large y = ?\]

OpenStudy (anonymous):

so it is 5

OpenStudy (johnweldon1993):

\[\large \frac{14}{2} \cancel={5} \]

OpenStudy (anonymous):

so it not 5

OpenStudy (johnweldon1993):

Well 14 divided by 2 is definitely not 5

OpenStudy (anonymous):

it is 7

OpenStudy (johnweldon1993):

Great!

OpenStudy (anonymous):

sorry i was look at to Questions at the same time